Determine the amount of money in a savings account at the end of 10 years, given an initial deposit of $5,500 and a 12 percent annual interest rate when interest is compounded

(a) annually,

(b) semiannually, and

(c) quarterly.

Include financial calculator steps, including the keys pressed on the calculator to solve each step of the question.
Answer #1

a.

PV = -5500

I/Y = 12

N = 10

PMT = 0

CPT -> FV

FV = 17082.17 Answer

b.

PV = -5500

I/Y = 12/2 = 6%

N = 10 *2 = 20

PMT = 0

CPT -> FV

FV = 17639.25 Answer

c.

PV = -5500

I/Y = 12/4 =3%

N = 10 *4 =40

PMT = 0

CPT -> FV

FV = 17941.21 Answer

Kindly do inform me in case you have any queries.